# Balanced rectangles over Sturmian words and minimal discrepancy intervals

### v1

@@ 5.3.  One-sided approximations
By our assumptions and by Lemma 5.5, we have $\langle n\alpha\rangle=\lVert n\alpha\rVert$ and $\langle n^{\prime}\alpha\rangle=\lVert n^{\prime}\alpha\rVert$. Thus, the inequality is the statement (7) in Lemma 5.1. ∎
@@
Let $L\geq 1$ (or $L\geq 0$, if $\alpha<1/2$). Let $\langle\cdot\rangle\in\{\langle\cdot\rangle_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{right}}\}$ be such that $\langle q_{L}\alpha\rangle<1/2$. Then
(5.9) $\langle q_{L}\alpha\rangle>\langle(q_{L}+q_{L+1})\alpha\rangle>\dots>\langle(q_{L}+(a_{L+2}-1)q_{L+1})\alpha\rangle>\langle q_{L+2}\alpha\rangle.$
@@
(5.10) $0<q<q_{L}+(a+1)q_{L+1}\quad\implies\quad\langle q\alpha\rangle\geq\langle(q_{L}+aq_{L+1})\alpha\rangle.$
@@
For the best approximation property (5.10), note that all $q$ in the range $0<q<q_{L+2}$ which are not of the shape $q_{L}+aq_{L+1}$ with $0\leq q\leq a_{L+2}$, by (5.6), must either have $k_{0}(q)=L+1$, or $k_{0}(q)=L$ and $b_{L}(q)>1$, or $k_{0}(q)\leq L-1$.
If $k_{0}(q)=L+1$, then Lemma 5.2 implies $\langle q\alpha\rangle>1/2>\langle q_{L}\alpha\rangle$.
If $k_{0}(q)\leq L-1$, then Lemma 5.6 implies $\langle q\alpha\rangle>\langle q_{L}\alpha\rangle$.
If $k_{0}(q)=L$ and $b_{L}(q)\geq 2$, then in the case $L\geq 2$ (or $L\geq 1$ if $\alpha<1/2)$, Lemma 5.7 implies $\langle q\alpha\rangle>\langle q_{L}\alpha\rangle$. In case $L=0$ (or $L=1$ if $\alpha>1/2$), this is easy to see as well.
Overall, $\langle q\alpha\rangle>\langle q_{L}\alpha\rangle$ holds for all $q$ that are not of the shape $q_{L}+aq_{L+1}$, and so the implication (5.10) follows from (5.9). ∎
@@ 5.4.  Minimising distances in certain ranges of integers
Assume that $k_{0}(n)=L$ with $L\geq 2$ (or $L\geq 1$ if $\alpha<1/2$). Set $\langle\cdot\rangle$ as in (5.8). Then
$\langle(n+q_{L+1})\alpha\rangle<\langle n\alpha\rangle.$
More generally, the inequality holds for all $n$ with $k_{0}(n)\leq L$ and $\langle n\alpha\rangle<1/2$, where $\langle\cdot\rangle$ is chosen such that $\langle q_{L}\alpha\rangle<1/2$.
@@
This follows from the fact that $\langle n\alpha\rangle=\lVert n\alpha\rVert>\lVert q_{L+1}\alpha\rVert$ (by Lemma 5.1(1)) and $\langle q_{L+1}\alpha\rangle=1-\lVert q_{L+1}\alpha\rVert$ (by Lemma 5.5, since $L$ and $L+1$ have distinct parities). ∎
@@
Assume that $k_{0}(n)=0$ and set $\langle\cdot\rangle$ as in (5.8). Then we have
$\langle(n-q_{1})\alpha\rangle<\langle n\alpha\rangle\quad\text{or}\quad\langle(n+q_{1})\alpha\rangle<\langle n\alpha\rangle.$
If $\alpha>1/2$, then the same holds for $k_{0}(n)=1$ and $\langle(n\pm q_{2})\alpha\rangle$.
@@
This follows from the fact that $\langle n\alpha\rangle=\lVert n\alpha\rVert>\lVert q_{1}\alpha\rVert$, by Lemma 5.1(1). ∎
@@
Assume $b_{0}(n)\geq 2$ and set $\langle\cdot\rangle$ as in (5.8). Then we have
$\langle(n-1)\alpha\rangle<\langle n\alpha\rangle\quad\text{or}\quad\langle(n+1)\alpha\rangle<\langle n\alpha\rangle.$
If $\alpha>1/2$, then the above statement holds with the assumption $b_{1}(n)\geq 2$.
@@
Let $\alpha<1/2$ and $b_{0}(n)\geq 2$. Then, similarly to the proof of Lemma 5.3, one can use (5.3) and (5.7) to show that $\left\{\alpha\right\}<\left\{n\alpha\right\}<1-\left\{\alpha\right\}$. This implies $\left\{(n-1)\alpha\right\}<\left\{n\alpha\right\}<\left\{(n+1)\alpha\right\}$, and so $\langle(n-1)\alpha\rangle_{\textnormal{left}}<\langle n\alpha\rangle_{\textnormal{left}}$ and $\langle(n+1)\alpha\rangle_{\textnormal{right}}<\langle n\alpha\rangle_{\textnormal{right}}$. This proves the lemma in the case that $\alpha<1/2$. For $\alpha>1/2$ the proof is analogous. ∎
@@
Assume that either $k_{0}(n)=L$ or $n=q_{L-1}+\sum_{k=L+2t}^{N}b_{k}q_{k}$ with $t\geq 0$ and $b_{L+2t}\neq 0$. Set $\langle\cdot\rangle$ as in (5.8). Then
(5.11) $\langle n\alpha\rangle=\min\{\langle x\alpha\rangle\colon n-q_{L}<x<n+q_{L}\}.$
@@
Assume first that $k_{0}(n)=L$. If $L=0$ (or $L=1$ if $\alpha>1/2$), then $q_{L}=1$, so the statement is trivial. Otherwise, by Lemma 5.1(2), we have $\lVert n\alpha\rVert<\lVert q_{L-1}\alpha\rVert$. On the other hand, by the best approximation property of convergents (5.2), we have $\lVert d\alpha\rVert\geq\lVert q_{L-1}\alpha\rVert$ for all $1\leq d<q_{L}$. Thus, for all $1\leq d<q_{L}$ with $\langle d\alpha\rangle<1/2$, we have $\langle(n+d)\alpha\rangle=\langle n\alpha\rangle+\lVert d\alpha\rVert>\langle n\alpha\rangle$ and $\langle(n-d)\alpha\rangle>1/2>\langle n\alpha\rangle$. Similarly, for all $1\leq d<q_{L}$ with $\langle d\alpha\rangle>1/2$, we have $\langle(n+d)\alpha\rangle>1/2>\langle n\alpha\rangle$ and $\langle(n-d)\alpha\rangle=\langle n\alpha\rangle+\lVert d\alpha\rVert>\langle n\alpha\rangle$. This implies (5.11).
Now assume that $n=q_{L-1}+\sum_{k=L+2t}^{N}b_{k}q_{k}$ with $t\geq 0$ and $b_{L+2t}\neq 0$. Then we know from Lemma 5.1(8) that again $\lVert n\alpha\rVert<\lVert q_{L-1}\alpha\rVert$, and so we can argue as in the previous case. ∎
@@
Assume that $k_{0}(n)\geq L$ with $L\geq 2$ (or $L\geq 1$ if $\alpha<1/2$). Set $\langle\cdot\rangle$ as in (5.8). Then
$\langle n\alpha\rangle=\min\{\langle x\alpha\rangle\colon n\leq x<n+q_{L-1}+q_{L}\}.$
@@
$\langle n\alpha\rangle=\min\{\langle x\alpha\rangle\colon n\leq x<n+q_{L-1}+2q_{L}\}.$
@@
Assume $k_{0}(n)=L$ with $L\geq 2$ (or $L\geq 1$, if $\alpha<1/2$) and $\langle n\alpha\rangle<1/2$. Then Lemma 5.1(2) implies $\langle n\alpha\rangle=\lVert n\alpha\rVert<\lVert q_{L-1}\alpha\rVert$. Let $\langle\cdot\rangle^{\prime}$ denote the distance to the nearest integer in the opposite direction, i.e., $\{\langle\cdot\rangle,\langle\cdot\rangle^{\prime}\}=\{\langle\cdot\rangle_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{right}}\}$. Then by Lemma 5.5 we have $\langle q_{L}\alpha\rangle<1/2$ and $\langle q_{L-1}\alpha\rangle^{\prime}<1/2$.
For the sake of contradiction, assume $\langle(n+d)\alpha\rangle<\langle n\alpha\rangle$ for some $1\leq d<q_{L-1}+q_{L}$. This is equivalent to $\langle(n+d)\alpha\rangle^{\prime}>\langle n\alpha\rangle^{\prime}$, and so
$\langle d\alpha\rangle^{\prime}=\langle(n+d)\alpha\rangle^{\prime}-\langle n\alpha\rangle^{\prime}<1-\langle n\alpha\rangle^{\prime}=\langle n\alpha\rangle<\lVert q_{L-1}\alpha\rVert=\langle q_{L-1}\alpha\rangle^{\prime}$
@@
Finally, assume that $k_{0}(n)=L$ and $b_{L}\leq a_{L+1}-1$. Then Lemma 5.1(6) says that in fact $\langle n\alpha\rangle<\lVert(q_{L-1}+q_{L})\alpha\rVert$, and we get the better bound by the same argument as before, this time using Lemma 5.8 with $a=1$. ∎
Next, we determine integers larger than $n$ which minimise $\langle x\alpha\rangle$ in certain ranges strictly above $n$.
@@
Assume that $\langle n\alpha\rangle<1/2$, and let
$n^{\prime}:=\min\{x>n\colon\langle x\alpha\rangle<\langle n\alpha\rangle\}.$
Set $u:=\min\{n^{\prime},n+q_{1}\}$ if $\alpha<1/2$, and $u:=\min\{n^{\prime},n+q_{2}\}$ if $\alpha>1/2$. Then
$\langle(n+1)\alpha\rangle=\min_{n<x<u}\langle x\alpha\rangle.$
Proof.
Assume first $\alpha<1/2$ and $\langle n\alpha\rangle<1/2$ and $n^{\prime}>n+1$ (if $n^{\prime}=n+1$, the statement is trivial). Then $\langle n\alpha\rangle<\langle(n+1)\alpha\rangle$. For the sake of contradiction, assume that there exists an integer $x$ with $n<x<u$ such that $\langle n\alpha\rangle<\langle x\alpha\rangle<\langle(n+1)\alpha\rangle$. Then we can write $x=n+d$ with some integer $1\leq d<q_{1}$, and we have $\langle n\alpha\rangle<\langle(n+d)\alpha\rangle<\langle(n+1)\alpha\rangle$. This implies $\langle d\alpha\rangle<\langle\alpha\rangle=\lVert\alpha\rVert$, which is impossible for $1\leq d<q_{1}$, by (5.2). The proof for $\alpha>1/2$ is analogous. ∎
Lemma 5.15.
Let $L\geq 1$ (or $L\geq 0$ if $\alpha<1/2$) and assume $\langle q_{L}\alpha\rangle<1/2$. Let $n\geq 1$ and
$n^{\prime}:=\min\{x>n\colon\langle x\alpha\rangle<\langle n\alpha\rangle\}.$
@@
$\langle(n+q_{L}+aq_{L+1})\alpha\rangle=\min_{n<x<u}\langle x\alpha\rangle.$
@@
First, note that the assumption $n+q_{L}+aq_{L+1}<n^{\prime}$ implies $\langle(n+q_{L}+aq_{L+1})\alpha\rangle>\langle n\alpha\rangle$. For the sake of contradiction, assume there exists an integer $x$ with $n<x<u$ such that $\langle x\alpha\rangle<\langle(n+q_{L}+aq_{L+1})\alpha\rangle$. Then, since $u\leq n^{\prime}$ and $u\leq n+q_{L}+(a+1)q_{L+1}$, we have in fact
(5.12) $\langle n\alpha\rangle<\langle x\alpha\rangle=\langle(n+d)\alpha\rangle<\langle(n+q_{L}+aq_{L+1})\alpha\rangle$
with $0<d<q_{L}+(a+1)q_{L+1}$. The inequality (5.12) implies $\langle d\alpha\rangle<\langle(q_{L}+aq_{L+1})\alpha\rangle$, contradicting Lemma 5.8. ∎
@@
We start with a simple lemma that follows pretty much directly from the properties of Ostrowski representations collected in Lemma 5.1.
@@
Let $2\leq n\leq q_{N}$. Assume that $k_{0}(n)=L$ with $L\geq 2$ (or $L\geq 1$ if $\alpha<1/2$), but that $n$ is not of the shape $n=q_{L}+\sum_{k=L+1+2t}^{N}b_{k}q_{k}$ with $t\geq 0$ and $b_{L+1+2t}>0$.
Then for $T\geq N+2$ we have that $q_{T}-n$ is of the shape $q_{T}-n=q_{L-1}+\sum_{k=L+2t}^{T-1}b_{k}q_{k}$.
@@
If $n$ is not of the shape $n=q_{L}+\sum_{k=L+1+2t}^{N}b_{k}q_{k}$, then Lemma 5.1(8, 6) implies that $\lVert q_{L}\alpha\rVert\leq\lVert n\alpha\rVert<\lVert q_{L-1}\alpha\rVert-\lVert q_{T}\alpha\rVert$. Thus, we have $\lVert(q_{T}-n)\alpha\rVert\geq\lVert q_{L}\alpha\rVert-\lVert q_{T}\alpha\rVert$. Now Lemma 5.1(5) implies that $k_{0}(q_{T}-n)\leq L$. The fact that $\left\{(q_{T}-n)\alpha\right\}<1/2\iff\left\{n\alpha\right\}>1/2$ implies via Lemma 5.2 that in fact $k_{0}(q_{T}-n)=L-1$. Finally, the fact that $\lVert(q_{T}-n)\alpha\rVert\leq\lVert n\alpha\rVert+\lVert q_{T}\alpha\rVert<\lVert q_{L-1}\alpha\rVert$ implies via Lemma 5.1(8) that $q_{T}-n$ is of the shape $q_{L-1}+\sum_{k=L+2t}^{T-1}b_{k}q_{k}$. ∎
In preparation for our second lemma, we introduce some more notation: For $n=\sum_{k=0}^{N}b_{k}q_{k}$ and $M\geq 0$ let us define the “small part of nn”, obtained by discarding terms with indices larger than $M$:
$n^{[\leq M]}:=\sum_{k=0}^{\min\{M,N\}}b_{k}q_{k}.$
@@
Similarly, we define the “large part of nn”, namely
$n^{[\geq M]}:=\sum_{k=M}^{N}b_{k}q_{k}.$
@@
Now we state and prove our second lemma regarding the shape of $n$ and $q_{T}-n$.
@@
Assume that $n^{[\leq M-1]}>0$ and $n^{[\geq M]}>0$. Then if $q_{T-2}\geq n$, we have
@@
Note that the assumption $n^{[\leq M-1]}>0$ implies that $M\geq 1$ (or $M\geq 2$ if $\alpha>1/2$). Moreover, assume $n\leq q_{N}$ and $T\geq N+2$.
@@
First, note that $0<n_{\textnormal{small}}^{\prime}<q_{M}$. (This follows from the assumption $n^{[\leq M-1]}>0$ and (5.13).) Now we want to show that $k_{0}(n_{\textnormal{large}}^{\prime})\geq M$ and $k_{0}(n_{\textnormal{large}}^{\prime})\equiv M+1\pmod{2}$.
Choose $\langle\cdot\rangle\in\{\langle\cdot\rangle_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{right}}\}$ so that $\langle q_{M+1}\alpha\rangle<1/2$. Then by the case assumption and Lemmas 5.1(5) and 5.2, we have
$\langle n^{[\geq M+1]}\alpha\rangle=\lVert n^{[\geq M+1]}\alpha\rVert<\lVert q_{M}\alpha\rVert-\lVert q_{N+2}\alpha\rVert\leq\lVert q_{M}\alpha\rVert-\lVert q_{T}\alpha\rVert.$
Since, moreover, $\lVert q_{T}\alpha\rVert<\langle n^{[\geq M+1]}\alpha\rangle$, we get that
$\langle(-q_{T}+n^{[\geq M+1]})\alpha\rangle<\lVert q_{M}\alpha\rVert=1-\langle q_{M}\alpha\rangle.$
@@
$1-\lVert q_{M}\alpha\rVert=\langle q_{M}\alpha\rangle$ $<\langle q_{M}\alpha\rangle+\langle(-q_{T}+n^{[\geq M+1]})\alpha\rangle$
$=\langle(-q_{T}+n^{[\geq M+1]}+q_{M})\alpha\rangle<1,$
@@
$\langle n_{\textnormal{large}}^{\prime}\alpha\rangle=\langle(q_{T}-n^{[\geq k_{M}(n)]}-q_{M})\alpha\rangle<\lVert q_{M}\alpha\rVert.$
The fact that $\lVert n_{\textnormal{large}}^{\prime}\alpha\rVert=\langle n_{\textnormal{large}}^{\prime}\alpha\rangle<\lVert q_{M}\alpha\rVert$ implies that $k_{0}(n_{\textnormal{large}}^{\prime})\leq M$. Moreover, since $\langle q_{M+1}\alpha\rangle<1/2$, Lemma 5.5 implies that $k_{0}(n_{\textnormal{large}}^{\prime})\equiv M+1\pmod{2}$.

### v2

@@ 5.3.  One-sided approximations
By our assumptions and by Lemma 5.5, we have $\langle n\alpha\rangle_{*}=\lVert n\alpha\rVert$ and $\langle n^{\prime}\alpha\rangle_{*}=\lVert n^{\prime}\alpha\rVert$. Thus, the inequality is the statement (7) in Lemma 5.1. ∎
@@
Let $\alpha\in(0,1/2)$ and $L\geq 0$. Let $\langle\cdot\rangle_{*}\in\{\langle\cdot\rangle_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{right}}\}$ be such that $\langle q_{L}\alpha\rangle_{*}<1/2$. Then
(5.9) $\langle q_{L}\alpha\rangle_{*}>\langle(q_{L}+q_{L+1})\alpha\rangle_{*}>\dots>\langle(q_{L}+(a_{L+2}-1)q_{L+1})\alpha\rangle_{*}>\langle q_{L+2}\alpha\rangle_{*}.$
@@
(5.10) $0<q<q_{L}+(a+1)q_{L+1}\quad\implies\quad\langle q\alpha\rangle_{*}\geq\langle(q_{L}+aq_{L+1})\alpha\rangle_{*}.$
@@
For the best approximation property (5.10), note that all $q$ in the range $0<q<q_{L+2}$ which are not of the shape $q_{L}+aq_{L+1}$ with $0\leq a\leq a_{L+2}-1$, by (5.6), must either have $k_{0}(q)=L+1$, or $k_{0}(q)=L$ and $b_{L}(q)\geq 2$, or $k_{0}(q)\leq L-1$.
If $k_{0}(q)=L+1$, then Lemma 5.2 implies $\langle q\alpha\rangle_{*}>1/2>\langle q_{L}\alpha\rangle_{*}$.
If $k_{0}(q)\leq L-1$, then Lemma 5.6 implies $\langle q\alpha\rangle_{*}>\langle q_{L}\alpha\rangle_{*}$.
If $k_{0}(q)=L$ and $b_{L}(q)\geq 2$, then in the case $L\geq 1$, Lemma 5.7 implies $\langle q\alpha\rangle_{*}>\langle q_{L}\alpha\rangle_{*}$. If $L=0$, this is easy to see as well.
Overall, $\langle q\alpha\rangle_{*}>\langle q_{L}\alpha\rangle_{*}$ holds for all $q$ that are not of the shape $q_{L}+aq_{L+1}$, and so the implication (5.10) follows from (5.9). ∎
@@ 5.4.  Minimising distances in certain ranges of integers
Let $\alpha\in(0,1/2)$ and assume that $k_{0}(n)=L$ with $L\geq 1$. Set $\langle\cdot\rangle_{*}$ as in (5.8). Then
$\langle(n+q_{L+1})\alpha\rangle_{*}<\langle n\alpha\rangle_{*}.$
More generally, the inequality holds for all $n$ with $k_{0}(n)\leq L$, as long as $\langle n\alpha\rangle_{*}<1/2$ and $\langle q_{L}\alpha\rangle_{*}<1/2$.
@@
This follows from the fact that $\langle n\alpha\rangle_{*}=\lVert n\alpha\rVert>\lVert q_{L+1}\alpha\rVert$ (by Lemma 5.1(1)) and $\langle q_{L+1}\alpha\rangle_{*}=1-\lVert q_{L+1}\alpha\rVert$ (by Lemma 5.5, since $L$ and $L+1$ have distinct parities). ∎
@@
Assume that $k_{0}(n)=0$ and set $\langle\cdot\rangle_{*}$ as in (5.8). Then we have
$\langle(n-q_{1})\alpha\rangle_{*}<\langle n\alpha\rangle_{*}\quad\text{or}\quad\langle(n+q_{1})\alpha\rangle_{*}<\langle n\alpha\rangle_{*}.$
@@
This follows from the fact that $\langle n\alpha\rangle_{*}=\lVert n\alpha\rVert>\lVert q_{1}\alpha\rVert$, by Lemma 5.1(1). ∎
@@
Assume $b_{0}(n)\geq 2$ and set $\langle\cdot\rangle_{*}$ as in (5.8). Then we have
$\langle(n-1)\alpha\rangle_{*}<\langle n\alpha\rangle_{*}\quad\text{or}\quad\langle(n+1)\alpha\rangle_{*}<\langle n\alpha\rangle_{*}.$
@@
Similarly to the proof of Lemma 5.3, one can use (5.3) and (5.7) to show that $\left\{\alpha\right\}<\left\{n\alpha\right\}<1-\left\{\alpha\right\}$. This implies $\left\{(n-1)\alpha\right\}<\left\{n\alpha\right\}<\left\{(n+1)\alpha\right\}$, and so $\langle(n-1)\alpha\rangle_{\textnormal{left}}<\langle n\alpha\rangle_{\textnormal{left}}$ and $\langle(n+1)\alpha\rangle_{\textnormal{right}}<\langle n\alpha\rangle_{\textnormal{right}}$. This proves the lemma. ∎
@@
Assume that either $k_{0}(n)\geq L$ or $n=q_{L-1}+\sum_{k=L+2t}^{N}b_{k}q_{k}$ with $t\geq 0$ and $b_{L+2t}\neq 0$. Set $\langle\cdot\rangle_{*}$ as in (5.8). Then
(5.11) $\langle n\alpha\rangle_{*}=\min\{\langle x\alpha\rangle_{*}\colon n-q_{L}<x<n+q_{L}\}.$
@@
Let $n$ be as in the statement of the lemma. If $L=0$, then $q_{L}=1$, so the statement is trivial. Otherwise, by Lemma 5.1(2, 8), we have $\lVert n\alpha\rVert<\lVert q_{L-1}\alpha\rVert$. On the other hand, by the best approximation property of convergents (5.2), we have $\lVert d\alpha\rVert\geq\lVert q_{L-1}\alpha\rVert$ for all $1\leq d<q_{L}$. Thus, for all $1\leq d<q_{L}$ with $\langle d\alpha\rangle_{*}<1/2$, we have $\langle(n+d)\alpha\rangle_{*}=\langle n\alpha\rangle_{*}+\langle d\alpha\rangle_{*}>\langle n\alpha\rangle_{*}$ and $\langle(n-d)\alpha\rangle_{*}>1/2>\langle n\alpha\rangle_{*}$. Similarly, for all $1\leq d<q_{L}$ with $\langle d\alpha\rangle_{*}>1/2$, we have $\langle(n+d)\alpha\rangle_{*}>1/2>\langle n\alpha\rangle_{*}$ and $\langle(n-d)\alpha\rangle_{*}=\langle n\alpha\rangle_{*}+\lVert d\alpha\rVert>\langle n\alpha\rangle_{*}$. This implies (5.11). ∎
@@
Let $\alpha\in(0,1)$ and assume that $k_{0}(n)\geq L$ with $L\geq 1$. Set $\langle\cdot\rangle_{*}$ as in (5.8). Then
$\langle n\alpha\rangle_{*}=\min\{\langle x\alpha\rangle_{*}\colon n\leq x<n+q_{L-1}+q_{L}\}.$
@@
$\langle n\alpha\rangle_{*}=\min\{\langle x\alpha\rangle_{*}\colon n\leq x<n+q_{L-1}+2q_{L}\}.$
@@
Assume first that $k_{0}(n)=L$ with $L\geq 1$ and $\langle n\alpha\rangle_{*}<1/2$. Then Lemma 5.1(2) implies $\langle n\alpha\rangle_{*}=\lVert n\alpha\rVert<\lVert q_{L-1}\alpha\rVert$. Let $\langle\cdot\rangle_{**}$ denote the distance to the nearest integer in the opposite direction, i.e., $\{\langle\cdot\rangle_{*},\langle\cdot\rangle_{**}\}=\{\langle\cdot\rangle_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{right}}\}$. Then by Lemma 5.5 we have $\langle q_{L}\alpha\rangle_{*}<1/2$ and $\langle q_{L-1}\alpha\rangle_{**}<1/2$.
For the sake of contradiction, assume $\langle(n+d)\alpha\rangle_{*}<\langle n\alpha\rangle_{*}$ for some $1\leq d<q_{L-1}+q_{L}$. This is equivalent to $\langle(n+d)\alpha\rangle_{**}>\langle n\alpha\rangle_{**}$, and so
$\langle d\alpha\rangle_{**}=\langle(n+d)\alpha\rangle_{**}-\langle n\alpha\rangle_{**}<1-\langle n\alpha\rangle_{**}=\langle n\alpha\rangle_{*}<\lVert q_{L-1}\alpha\rVert=\langle q_{L-1}\alpha\rangle_{**}$
@@
Finally, assume that $k_{0}(n)=L$ and $b_{L}\leq a_{L+1}-1$. Then Lemma 5.1(6) says that in fact $\langle n\alpha\rangle_{*}<\lVert(q_{L-1}+q_{L})\alpha\rVert$, and we get the better bound by the same argument as before, this time using Lemma 5.8 with $a=1$. ∎
Next, we determine integers larger than $n$ which minimise $\langle x\alpha\rangle_{*}$ in certain ranges strictly above $n$.
@@
Let $\alpha\in(0,1)$ and $L\geq 0$, and assume $\langle q_{L}\alpha\rangle_{*}<1/2$. Let $n\geq 1$ and
$n^{\prime}:=\min\{x>n\colon\langle x\alpha\rangle_{*}<\langle n\alpha\rangle_{*}\}.$
@@
$\langle(n+q_{L}+aq_{L+1})\alpha\rangle_{*}=\min_{n<x<u}\langle x\alpha\rangle_{*}.$
@@
First, note that the assumption $n+q_{L}+aq_{L+1}<n^{\prime}$ implies $\langle(n+q_{L}+aq_{L+1})\alpha\rangle_{*}>\langle n\alpha\rangle_{*}$. For the sake of contradiction, assume there exists an integer $x$ with $n<x<u$ such that $\langle x\alpha\rangle_{*}<\langle(n+q_{L}+aq_{L+1})\alpha\rangle_{*}$. Then, since $u\leq n^{\prime}$ and $u\leq n+q_{L}+(a+1)q_{L+1}$, we have in fact
(5.12) $\langle n\alpha\rangle_{*}<\langle x\alpha\rangle_{*}=\langle(n+d)\alpha\rangle_{*}<\langle(n+q_{L}+aq_{L+1})\alpha\rangle_{*}$
with $0<d<q_{L}+(a+1)q_{L+1}$. But (5.12) implies $\langle d\alpha\rangle_{*}<\langle(q_{L}+aq_{L+1})\alpha\rangle_{*}$, contradicting Lemma 5.8. ∎
@@
Let $\alpha\in(0,1/2)$ and assume that $k_{0}(n)=L$ with $L\geq 1$ and $q_{L}<n<q_{N+1}$, but that $n$ is not of the shape $n=q_{L}+\sum_{k=L+1+2t}^{N}b_{k}q_{k}$ with $t\geq 0$ and $b_{L+1+2t}>0$.
Then for $T\geq N+2$ we have that $q_{T}-n$ is of the shape $q_{T}-n=q_{L-1}+\sum_{k=L+2t}^{T-1}b^{\prime}_{k}q_{k}$ with $t\geq 0$ and $b^{\prime}_{L+2t}>0$.
@@
Let $n$ be as in the statement of the lemma. Then by Lemma 5.1(9) we have $\lVert n\alpha\rVert\geq\lVert q_{L}\alpha\rVert+\lVert q_{N}\rVert>\lVert q_{L}\alpha\rVert+\lVert q_{T}\alpha\rVert$, and so $\lVert(q_{T}-n)\alpha\rVert>\lVert q_{L}\alpha\rVert$. Now Lemma 5.1(2) implies that $k_{0}(q_{T}-n)\leq L$. Note that $\left\{(q_{T}-n)\alpha\right\}<1/2\iff\left\{n\alpha\right\}>1/2$. Thus, by Lemma 5.2, $k_{0}(n)$ and $k_{0}(q_{T}-n)$ must have distinct parities, and so in fact $k_{0}(q_{T}-n)\leq L-1$.
Finally, since $k_{0}(n)=L$, Lemma 5.1(5) tells us that $\lVert n\alpha\rVert<\lVert q_{L-1}\alpha\rVert-\lVert q_{N+2}\rVert\leq\lVert q_{L-1}\alpha\rVert-\lVert q_{T}\rVert$, and so $\lVert(q_{T}-n)\alpha\rVert<\lVert q_{L-1}\alpha\rVert$. Thus, we get from Lemma 5.1(1, 8) that $q_{T}-n$ is of the shape $q_{L-1}+\sum_{k=L+2t}^{T-1}b^{\prime}_{k}q_{k}$ with $t\geq 0$ and $b^{\prime}_{L+2t}>0$. ∎
In preparation for our second lemma, we introduce some more notation: For $M\geq 0$ let us define the “small part of nn”, obtained by discarding terms with indices larger than $M$:
$n^{[\leq M]}:=\sum_{k=0}^{M}b_{k}(n)q_{k}.$
@@
Similarly, we define the “large part of nn”:
$n^{[\geq M]}:=\sum_{k=M}^{\infty}b_{k}(n)q_{k}.$
@@
Assume that $n^{[\leq M-1]}>0$ and $n^{[\geq M]}>0$. Then if $q_{T-1}>n$, we have
@@
Note that the assumption $n^{[\leq M-1]}>0$ implies that $M\geq 1$. Moreover, assume $n<q_{N+1}$ and $T\geq N+2$.
@@
First, note that $0<n_{\textnormal{small}}^{\prime}<q_{M}$. (This follows from the assumption $n^{[\leq M-1]}>0$ and (5.13).) Our goal is to show that $k_{0}(n_{\textnormal{large}}^{\prime})\geq M$ and $k_{0}(n_{\textnormal{large}}^{\prime})\equiv M+1\pmod{2}$.
Choose $\langle\cdot\rangle_{*}\in\{\langle\cdot\rangle_{\textnormal{left}},\langle\cdot\rangle_{\textnormal{right}}\}$ so that $\langle q_{M+1}\alpha\rangle_{*}<1/2$. Then by the case assumption and Lemmas 5.1(5) and 5.2, we have
$\langle n^{[\geq M+1]}\alpha\rangle_{*}=\lVert n^{[\geq M+1]}\alpha\rVert<\lVert q_{M}\alpha\rVert-\lVert q_{N+2}\alpha\rVert\leq\lVert q_{M}\alpha\rVert-\lVert q_{T}\alpha\rVert.$
Since, moreover, $\lVert q_{T}\alpha\rVert<\langle n^{[\geq M+1]}\alpha\rangle_{*}$, we get that
$\langle(-q_{T}+n^{[\geq M+1]})\alpha\rangle_{*}<\lVert q_{M}\alpha\rVert=1-\langle q_{M}\alpha\rangle_{*}.$
@@
$1-\lVert q_{M}\alpha\rVert=\langle q_{M}\alpha\rangle_{*}$ $<\langle q_{M}\alpha\rangle_{*}+\langle(-q_{T}+n^{[\geq M+1]})\alpha\rangle_{*}$
$=\langle(-q_{T}+n^{[\geq M+1]}+q_{M})\alpha\rangle_{*}<1,$
@@
$\langle n_{\textnormal{large}}^{\prime}\alpha\rangle_{*}=\langle(q_{T}-n^{[\geq M+1]}-q_{M})\alpha\rangle_{*}<\lVert q_{M}\alpha\rVert.$
The fact that $\lVert n_{\textnormal{large}}^{\prime}\alpha\rVert=\langle n_{\textnormal{large}}^{\prime}\alpha\rangle_{*}<\lVert q_{M}\alpha\rVert$ implies that $k_{0}(n_{\textnormal{large}}^{\prime})\geq M$. Moreover, since $\langle q_{M+1}\alpha\rangle_{*}<1/2$, Lemma 5.5 implies that $k_{0}(n_{\textnormal{large}}^{\prime})\equiv M+1\pmod{2}$.